# Naturalistic Computational Cognitive Science Towards generalizable models and theories that capture the full range of natural behavior

### v2

@@ 5.1 Frictionless reproducibility: a superpower
R3. Adopt the “single file” philosophy. Ideally, key workflows (e.g. model or evaluation definitions) should be defined in a single file. This has emerged as a powerful force for reproducibility in ML. While this approach contradicts traditional software engineering principles of modularity and abstraction, it offers distinct advantages for research code: (1) complete understanding of an important component (e.g. a modell) requires reading only one file, (2) minimal abstraction layers and dependencies, and (3) easy adaptation of prior methods or their components, by simply copying this single file into another codebase. Removing friction from future reuse also benefits us—as Table 1 shows, single-file methods get many citations.
5.2 Developing models with an eye towards generalizability
Task-performing computational models have been a goal of cognitive science since at least  Newell, (1973). This sentiment persists today (Kriegeskorte and Douglas,, 2018; Almaatouq et al.,, 2024) and is exemplified in conferences that explicitly seek to integrate cognitive science and neuroscience with AI  (Naselaris et al.,, 2018). Cognitive scientists are well aware that relying on a single paradigm to study an effect may reduce the generalizability of conclusions (Holzmeister et al.,, 2024; Yarkoni,, 2022). Ideally, the models we develop generalize beyond the setting they were designed for. This requires two things. First, that the model is generalizable, i.e. can learn from, operate over, and perform new tasks. Second, that the behavioral (and potentially neural) predictions made by our models capture data on these new tasks. Clearly, the first is a requirement for the second. Here, we detail suggestions from ML for how cognitive science can develop more generalizable models.
Recommendations for cognitive science
R4. Develop models with simplified tasks, but ensure they work with increasingly naturalistic tasks. The frame problem (Shanahan,, 2004) and the “simulation is doomed to succeed” critique (Grim et al.,, 2013) highlight a fundamental challenge in model development: the evaluation tasks we design often fail to capture the full complexity of the phenomenon we care about. mchine learning addresses this by developing models that must succeed in both simplified and increasingly complex settings. The history of RL shows one such progression, focused on increasing naturalism: initial environments were grid-worlds (Sutton and Barto,, 2018), later 2D video games (Bellemare et al.,, 2013), then virtual 3D homes (Szot et al.,, 2021; Li et al., 2021b, ), and now open-world environments (Fan et al.,, 2022; Matthews et al.,, 2024; Hughes et al.,, 2024). Despite this increased emphasis on naturalistic settings, there is still emphasis on the value of simpler settings for studying specific aspects of models (Osband et al.,, 2019; Obando Ceron et al.,, 2024). Indeed, the strategy is generally to first develop your model in simplified settings, but then to verify that your model continues to “work” with more naturalistic settings. Collectively, this has allowed RL researchers to advance both theoretical (Dadashi et al.,, 2019; Grimm et al.,, 2020; Lyle et al.,, 2022) and empirical (Obando Ceron et al.,, 2024; Farebrother et al.,, 2024) understanding of how their models work, but also develop performant models for many“complex” settings, including beating world experts in Go (Silver et al.,, 2017) and controlling real robots (Levine et al.,, 2016; Cheng et al.,, 2024).
R5. Develop models that work on many task. The ultimate goal of ML is generalization (Highleyman and Kamentsky,, 1959; Hardt and Recht,, 2022; Recht,, 2024). Initially, this meant models that generalize to new data. Now, it means both models and learning procedures that generalize to new settings. Thus, the practice of evaluating models across diverse datasets to showcase their generality has become foundational in ML. Early examples include “dropout” (Srivastava et al.,, 2014), which demonstrated improvements across 6 image and speech datasets, and MoCo (He et al.,, 2020), which validated its approach on 9 vision datasets. Another notable example is the seminal “Deep Q-Network” paper (Mnih et al.,, 2015), which showed that a single neural network architecture with fixed hyperparameters [^2] E.g., the number of layers or the learning rate. could master 50 different Atari games. The field has since evolved to organize environments by the cognitive capabilities they test (e.g., exploration, generalization, manipulation) (Patterson et al.,, 2023), with researchers developing models on some task sets and validating on entirely different ones to ensure robust generalization (Machado et al.,, 2018).
We recommend cognitive scientists also test the models that support their theories more broadly. Importantly, we should evaluate models on different tasks from those we use to develop them (Machado et al.,, 2018). This further helps prevent our model (and thereby theory) from “overfitting” to a particular paradigm.
R6. Iterate between (1) model-design (2) task design where the other is fixed. One crucial lesson from ML research is the importance of separating model development from task design. Patterson et al., (2023) strongly caution against simultaneously developing both the problem setting (datasets or tasks) and the solution method, as researchers may inadvertently design evaluations that favor their proposed solution. Historical examples illustrate this risk. When developing Q-learning algorithms, researchers modified the classic pendulum control problem by adding episode cutoffs and random state resets (Machado et al.,, 2018). These modifications made Q-learning more tractable by implicitly improving exploration, but thereby obscured fundamental limitations in Q-learning’s exploration compared to policy gradient methods. Maintaining a strict separation between task design and model development enforces a clear delineation between hypothesis formation and testing, which is important for reducing experimenter bias. This is especially important because researchers tend to find “positive” results (Scheel et al.,, 2021; Sarafoglou et al.,, 2022), a bias exacerbated by an increasingly competitive academic landscape (Lee,, 2014; Reithmeier et al.,, 2019; Woolston,, 2021). Much of ML’s progress can be attributed to researchers evaluating their models on independently developed datasets and tasks—a practice that both simplifies the research process and promotes scientific rigor.
Thus, we recommend that cognitive science also develop models on tasks/stimuli developed by others. This may seem to contradict science, where researchers craft experiments to test their specific hypotheses. The data from prior work was for a different hypothesis! We recommend that the field develop phenomena-oriented benchmarks that aggregate stimuli and tasks designed to test specific theoretical predictions for broad phenomena. This approach would allow researchers to leverage existing experimental paradigms while maintaining scientific rigor through independent validation. We detail this further in the next section.
5.3 Benchmarks: catalysts for progress and innovation
Another challenge for cognitive science is fragmentation. Newell’s seminal critique (Newell,, 1973)—that piece-wise investigation of brain components would not yield a cohesive understanding—remains relevant forty years later (Schrimpf et al.,, 2020; DiCarlo et al.,, 2023). Current experimental paradigms often produce disconnected, incommensurate findings that resist integration across studies, even within a domain (Almaatouq et al.,, 2024). This fragmentation manifests not only in methodology but also in how researchers distribute their attention across different problems, thereby diluting collective progress.
The history of machine learning offers an instructive parallel. Initially, researchers worked on disparate, self-defined problems, leading to diffuse progress (Recht,, 2024). The introduction of shared benchmarks drove a crucial shift: researchers began focusing on standardized problems that others had identified as important. Putting aside whether this is good or bad, it has been effective. Indeed, this is the third pillar of frictionless reproducibility which Donoho, (2024) argues has been a major catalyst for progress in ML: competitive challenges. We argue that cognitive science can similarly stand to benefit from benchmarks.
Recommendations for cognitive science
R7. Embrace the leaderboard principle. A common concern with benchmarks is that researchers will “overfit” to test sets, which theoretically should underestimate the true test error (Hastie et al.,, 2009). Yet despite widespread iteration on benchmarks in machine learning the field has produced models with strong performance on held-out data. This apparent paradox is partially explained by a phenomenon known as the “leaderboard principle” (Hardt and Recht,, 2022). Researchers typically only publish and build upon models that demonstrate substantial improvements over prior results, rather than minor variations. This selective pressure towards meaningful advances effectively constrains the degree of adaptation to test sets, as researchers focus on substantial improvements rather than exhaustively exploring the test set’s peculiarities. Nevertheless, a risk of overfitting to data remains. We argue that cognitive science can avoid this through the creation of dynamic phenomena-oriented benchmarks, where researchers can contribute new challenges that identify a missing component of a phenomenon of interest. We detail this below.
R8. Develop dynamic benchmarks to iteratively capture important phenomena. ML benchmarks are not collected to test specific hypotheses (Hardt and Recht,, 2022) and so may seem unscientific. However, cognitive science can embrace “scientific” benchmarks oriented toward testing specific hypotheses. We use the story of “Brain-Score” as an illustrative example of initial steps in this direction.
Brain-Score (Schrimpf et al.,, 2020) began as a benchmark collecting behavioral and neural data from humans and monkeys on object recognition tasks. One common critique of Brain-Score is that the behavioral analysis essentially measures whether models and humans found the same stimuli easy or challenging to identify. However, it did not test hypotheses about how humans identify objects, nor show that models identified objects in this same way. For example, Bowers et al., (2023) noted that matching human predictions is insufficient to identify a good model of human object recognition. The authors reference 9 experiments from the psychophysics literature which presented stimuli generated to test specific hypotheses, and showed that deep neural networks failed to capture human behavior across these. Examples include stimuli testing the hypothesis that object recognition was based on local shape features (Baker et al.,, 2020), stimuli testing the hypothesis that human object recognition is defined by shape more than by texture (Geirhos et al.,, 2018), and stimuli testing the hypothesis that pairs of objects were judged to be similar based on shape contours (Puebla and Bowers,, 2022). In response, DiCarlo et al., (2023) added many of these stimulus sets to Brain-Score so that when future models are tested on whether they capture human behavior across all of these stimulus sets.
This story illustrates what an exciting strand of computational cognitive science research could look like: one where researchers develop models and experiments that easily integrate with prior findings on a phenomenon thanks to phenomena-oriented benchmarks, frictionless reproducibility (§5.1), and generalizable models (§5.2). By combining these, we achieve a robust foundation for building theories that span the full space of natural behavior associated with a phenomenon.
We are not the first to identify that cognitive science is fragmented and to suggest benchmarks that aggregate results from the literature as a solution. Indeed, researchers spanning memory, timing, confidence and more have already begun to do so (MacWhinney and Snow,, 1985; Oberauer et al.,, 2018; Aydoğan et al.,, 2024; Rahnev et al.,, 2020; Frank et al.,, 2017). Here, we both encourage more subdisciplines within cognitive science to follow suite but also emphasize the importance of frictionless reproducibility and building generalizable models. While prior work has begun the development of benchmarks, documentation for downloading and using the data is sometimes lacking; download access is not always authentication-free; and maybe most importantly, if one wants to add new data, there isn’t an easy way to re-run all previous models and ensure that findings are commensurate. Critically, while prior benchmarks have focused on aggregating experimental results from “simplified” experiments, we advocate for dynamic benchmarks whose development is oriented around explaining data from increasingly naturalistic experiments (i.e. experiments with either parametrically generated naturalistic stimuli or, eventually, stimuli taken from real-world conditions).
@@ 6 Building from naturalistic experiments to cognitive theories
Figure 7: Overview of how we can develop theories with potentially opaque models and unnatural manipulations of natural data. (Panel C bottom figure is reproduced from Saxe et al., 2019.)
A primary goal of cognitive science is understanding. By instantiating our theories in a model, we are forced to make our theories more precise by concretizing the ambiguous details of the mapping from a verbal hypothesis to its implementation (Guest and Martin,, 2021).
From this perspective, building computational models that perform the task in as naturalistic a way as possible, across as wide a variety of settings as possible, imposes much stronger constraints on our theories than an abstracted model at a higher-level (cf. Cao and Yamins, 2024b, ). These added constraints are important; otherwise, high-level theories that do not directly engage with the details of the implementation can be so unconstrained as to lack explanatory value (cf. Erev and Greiner,, 2015; Jones and Love,, 2011; Rahnev and Denison,, 2018; Andrews,, 2021), can be intractable for real problems (Van Rooij,, 2008), or can elide important details (see above).
Yet moving towards naturalistic settings, and the complex models they require, alters the way that we need to approach experimentation and theory building. Here, we discuss the consequences for experiment design (§6.1-6.2), the role and interpretation of models (6.3), and theory building (§6.4).
@@
Figure 8: Testing hypotheses by parametrically manipulating naturalistic data. In the natural data distribution (blue), shape and texture are highly correlated—cats usually have fur, and elephants usually have wrinkles. This confounding makes it difficult to discern how these variables independently contribute to processing in neural systems or models. One approach to deconfounding these variables is to use a narrow distribution of highly-simplified artificial stimuli (light green) that isolate the variables of interest. However, we advocate for (additionally) creating broader distributions of stimuli that preserve more of the breadth of the natural data distribution, while deconfounding or manipulating features parametrically in order to test hypotheses—as in the unnatural combinations of natural shapes and textures created by Geirhos et al., (2018).
@@
Developing and testing models on truly naturalistic data introduces new issues that are absent in simple artificial settings. Naturalistic data distributions may be too “easy”—e.g. they may only rarely include edge-cases that test key capabilities (Zhang et al.,, 2023)—thus, testing average performance over naturalistic data may disguise important failures. Relatedly, in naturalistic data, features may be confounded, which can prevent readily disentangling the solutions a model uses. For example, models trained on ImageNet (Deng et al.,, 2009) rely on features like texture more than humans do (Geirhos et al.,, 2018), but still perform similarly on ImageNet. These issues limit our ability to achieve deep understanding from experimenting with only the distribution of naturalistic data (Rust and Movshon,, 2005).
@@
However, we believe that these challenges are surmountable by parametrically manipulating rich, naturalistic data (Fig. 8)—preserving or enhancing its richness, while more carefully manipulating theoretically-motivated constructs (cf. Jain et al.,, 2024). For example, Geirhos et al., (2018) engineered datasets that combine natural shapes and textures in unnatural ways. To do so, the authors exploited the contemporary ML technique of iterative style transfer Gatys et al., (2016)—combining the features of different images at different spatial scales—thus illustrating how progress in AI models unlocks new experimental approaches. These conflicting shape-texture datasets helped to inspire similar experiments on humans (Jagadeesh and Gardner, 2022b, ; Jagadeesh and Gardner, 2022a, ); surprisingly, those studies found that human visual cortex likewise appears to use textural rather than shape-focused representations, and shape-driven behavior is thus mediated by downstream readout processes. Thus, manipulating features to test models can lead to new insights into brain function.
@@
A growing range of studies similarly systematically manipulate natural stimuli, while preserving as much of the natural variation as possible—whether introducing new parameters like stimulus modality manipulation (e.g. listening vs. reading; Deniz et al.,, 2019), inserting challenging syntactic structures in natural stories (Futrell et al.,, 2021), or systematically selecting stimuli that drive particular responses while preserving variety (e.g. Tuckute et al.,, 2024; Hosseini et al.,, 2023). Theory-driven, controlled experiments can be performed in more complex, naturalistic settings.
@@
Another direction for parametrically manipulating naturalistic data comes from using video games to study sequential decision-making and higher-level cognition (Allen et al.,, 2024). While games are artificial, different parameters such as the size of a world, the number of agents, the length of tasks, can be parametrically manipulated to reflect values more inline with those found in the real-world. As such, simulated environments used in ML offer valuable settings for cognitive experimentation. For example, Craftax (Matthews et al.,, 2024), a 2D Minecraft environment, offers parametrically-generated open-world environments containing complex goal hierarchies, which can be used for studying topics like exploration or planning. Other environments offer opportunities to study long-horizon generalization (Nikulin et al.,, 2024), coordination (Rutherford et al.,, 2024), physical reasoning (Matthews et al.,, 2025), etc. Because these environments are parametrically generated it is easy to create highly-variable-yet-controlled manipulations, e.g. many environments that systematically vary planning difficulty. Critically, a new tool called NiceWebRL [^3] https://github.com/KempnerInstitute/nicewebrl, enables easily running web-browser human experiments using the environemnts above and the many other ML environments written using Jax (Bradbury et al.,, 2018). Thus, there is a growing ecosystem of tools for building complex naturalistic experimental paradigms, and evaluating humans and models with them.
@@ 6.2 Incorporating naturalistic variation to enhance generalizability
By testing hypotheses in settings where other aspects of the task and data generating process are sampled broadly from the naturalistic distribution, we address some of the challenges of generalizability that arise from testing on a much narrower distribution than the written hypothesis implies (cf. Yarkoni,, 2022). In particular, testing our hypotheses across varied settings—“design[ing] with variability in mind” (Yarkoni,, 2022)—increases the likelihood of identifying robust effects that will generalize (cf. Baribault et al.,, 2018). In the language of regression, varying as many other data generating factors as possible brings us closer to estimating the “main effect” of an experimentally-manipulated construct, rather than estimating the “simple effect” of the manipulation in the single setting we have tested.
For example, it is a long-standing principle of experimental design to test a hypothesis with multiple stimuli, and statistically quantify the stimulus effect to estimate generalizability (e.g. Clark,, 1973). However, many other experiment features—even exact task formulations such as multi-arm bandit tasks—are often preserved within a paper, or even across an entire literature. Just like testing a hypothesis with a single stimulus, this limited variation poses challenges for generalizability (Eckstein et al.,, 2022)—especially because the way that researchers operationalize a theory can substantially affect their conclusions (e.g. Holzmeister et al.,, 2024; Schweinsberg et al.,, 2021; Strickland and Suben,, 2012). While naturalistic variation cannot resolve this issue in full, we believe that incorporating a broader range of the settings, stimuli, and tasks within which we expect a theory to hold, will help us to develop more generalizable theories.
Incorporating variation can also help us to revise the constructs underlying our theories. For example, Eisenberg et al., (2019) study the putative construct of self-regulation by simultaneously testing many existing self-regulation measures. The authors find that these measures do not form a unitary construct. Rather, task- and survey-based measures each form distinct, multi-factorial constructs, with individual tasks loading primarily onto a subset of factors. That is, two different studies that each use only a single measure for self-regulation may be measuring entirely different constructs. By contrast, by substantially varying the tasks we use to measure a theoretical component like self-regulation, we can more fully disentangle the underlying constructs—and thereby improve the extent to which our theories generalize.

### v4

@@ 5.3 Frictionless reproducibility: a superpower
R8. Code should be easy to “fork” and iterate on. There should be minimal friction for new researchers to take your existing code and try their ideas on it. One particularly promising strategy is the “single file” philosophy. Ideally, key workflows (e.g. model or evaluation definitions) should be defined in a single file. This has emerged as a powerful force for reproducibility in ML. While this approach contradicts traditional software engineering principles of modularity and abstraction, it offers distinct advantages for research code: (1) complete understanding of an important component (e.g. a model) requires reading only one file, (2) minimal abstraction layers and dependencies, and (3) easy adaptation of prior methods or their components, by simply copying this single file into another codebase. Removing friction from future reuse may also benefit us, as more people can build on and cite our work.
@@ 6 Building from naturalistic experiments to cognitive theories
Figure 7: Overview of how we can develop theories with potentially opaque models and unnatural manipulations of natural data: building task-performing models that can reproduce behavior across naturalistic and unnatural stimuli, while tightly coupling them to reductive explanations. (Panel C top figure is reproduced from Saxe et al. 2019; panel A images reproduced from Geirhos et al. (2018).
A primary goal of cognitive science is understanding. By instantiating our theories in a model, we are forced to make our theories precise by concretizing the ambiguous details of the mapping from a verbal hypothesis to its implementation (Guest and Martin, 2021).
From this perspective, building computational models that perform the task in as naturalistic a way as possible, across as wide a variety of settings as possible, imposes much stronger constraints on our theories than an abstracted model at a higher level (cf. Cao and Yamins, 2024b). These added constraints are important; otherwise, high-level theories that do not directly engage with the details of the implementation can be so unconstrained as to lack explanatory value (cf. Erev and Greiner, 2015; Jones and Love, 2011; Rahnev and Denison, 2018; Andrews, 2021), can be intractable for real problems (Van Rooij, 2008), or can elide important details (see above).
Yet moving towards naturalistic settings, and the complex models they require, alters the way that we need to approach experimentation and theory building. Here, we discuss the consequences for experiment design (§6.1-6.2), the role and interpretation of models (§6.3), and theory building (§6.4).
@@
Figure 8: Testing hypotheses by parametrically manipulating naturalistic data. In the natural data distribution (blue), shape and texture are highly correlated—cats usually have fur, and elephants usually have wrinkles. This confounding makes it difficult to discern how these variables independently contribute to processing in neural systems or models. One approach to deconfounding these variables is to use a narrow distribution of highly-simplified artificial stimuli (light green) that isolate the variables of interest. However, we advocate for (additionally) creating broader distributions of stimuli that preserve more of the breadth of the natural data distribution, while deconfounding or manipulating features parametrically in order to test hypotheses—as in the unnatural combinations of natural shapes and textures created by Geirhos et al. (2018).
@@
Developing and testing models on truly naturalistic data introduces new issues compared to simple artificial settings. Naturalistic data distributions may be too “easy”—e.g. they may only rarely include edge-cases that test key capabilities (Zhang et al., 2023)—thus, testing average performance over naturalistic data may disguise important failures. Relatedly, in naturalistic data, features may be confounded, which can prevent readily disentangling the solutions a model uses. For example, models trained on ImageNet (Deng et al., 2009) rely on features like texture more than humans do (Geirhos et al., 2018), but still perform similarly on ImageNet. These issues limit our ability to achieve deep understanding when experimenting with only the naturalistic distribution (Rust and Movshon, 2005).
@@
However, these challenges are surmountable by parametrically manipulating rich, naturalistic data (Fig. 8)—preserving or enhancing its richness, while still carefully manipulating theoretically-motivated constructs (cf. Jain et al., 2024). For example, Geirhos et al. (2018) engineered datasets that combine natural shapes and textures in unnatural ways. To do so, the authors exploited the contemporary ML technique of iterative style transfer (Gatys et al., 2016)—combining features of different images at different spatial scales—thus illustrating how progress in AI models unlocks new experimental approaches. These conflicting shape-texture datasets helped to inspire similar experiments on humans (Jagadeesh and Gardner, 2022a, b); surprisingly, those studies found that human visual cortex likewise appears to use textural rather than shape-focused representations. Thus, shape-driven behavior is mediated by downstream readout processes. Thus, manipulating naturalistic features to test models can lead to new insights into brain function.
@@
Many studies similarly systematically manipulate natural stimuli, while preserving as much natural variation as possible—whether introducing new parameters like modality manipulation (e.g. listening vs. reading; Deniz et al., 2019), inserting challenging syntactic structures in natural stories (Futrell et al., 2021), or systematically selecting stimuli that drive particular responses while preserving variety (e.g. Tuckute et al., 2024; Hosseini et al., 2023). Theory-driven, controlled experimental manipulations can be performed even in more complex, naturalistic settings.
@@
Another direction for parametrically manipulating naturalistic data comes from using video games to study sequential decision-making and higher-level cognition (Allen et al., 2024). Games combine rich environments with the ability to parametrically manipulate features to reflect naturalistic values, while introducing experimental manipulation. For example Carvalho et al. (2025) tested human goal preplay using Craftax (Matthews et al., 2024)—a 2D Minecraft environment that offers parametrically-generated open-world environments containing complex goal hierarchies. By sampling complex maps, while constraining possible paths and goal locations, the authors created varied naturalistic experiments that could nevertheless cleanly disentangle different models of goal generalization. For researchers who want to build on games, there is a growing ecosystem of tools for building complex naturalistic experimental paradigms, and evaluating humans and models with them— such as NiceWebRL [^4] https://github.com/KempnerInstitute/nicewebrl, which enables web-browser human experiments using ML environments.
@@ 6.2 Incorporating naturalistic variation to enhance generalizability
By testing hypotheses in settings where other aspects of the task and data generating process are sampled broadly from the naturalistic distribution, we address some of the generalizability challenges that arise from testing on a much narrower distribution than the written hypothesis implies (cf. Yarkoni, 2022); “design[ing] with variability in mind” (Yarkoni, 2022) increases the likelihood of identifying robust effects that will generalize (cf. Baribault et al., 2018). In the language of regression, varying as many other data generating factors as possible brings us closer to estimating the “main effect” of an experimentally-manipulated construct, rather than estimating the “simple effect” of the manipulation in the single setting we have tested.
For example, it is a long-standing principle of experimental design to test a hypothesis with multiple stimuli, and statistically quantify the stimulus effect to estimate generalizability (e.g. Clark, 1973). However, many other experiment features—even exact task formulations such as multi-arm bandit tasks—are often preserved within a paper, or even across an entire literature. Just like testing a hypothesis with a single stimulus, this limited variation poses challenges for generalizability (Eckstein et al., 2022)—especially because the way that researchers operationalize a theory can substantially affect their conclusions (e.g. Holzmeister et al., 2024; Schweinsberg et al., 2021; Strickland and Suben, 2012). While naturalistic variation cannot resolve this issue in full, we believe that incorporating a broader range of the settings, stimuli, and tasks within which we expect a theory to hold, will help us to develop more generalizable theories.
Incorporating variation can also help us to revise the constructs underlying our theories. For example, Eisenberg et al. (2019) study the putative construct of self-regulation by simultaneously testing many existing self-regulation measures. The authors find that these measures do not form a unitary construct. Rather, task- and survey-based measures each form distinct, multi-factorial constructs, with individual tasks loading primarily onto a subset of factors. That is, two different studies that each use only a single measure for self-regulation may be measuring entirely different constructs. By contrast, by substantially varying the tasks we use to measure a theoretical component like self-regulation, we can more fully disentangle the underlying constructs—and thereby improve the extent to which our theories generalize. Likewise, Peterson et al. (2021) collect a large dataset of risky choice judgments, and use it to explore how human choices interpolate between different prior decision-making theories depending on the problem context.